A canoeist paddles 6 miles downstream in 40 minutes and returns in 3 hours paddling at the same rate. How fast does he go in still water?
Answer by lwsshak3(11628)

A canoeist paddles 6 miles downstream in 40 minutes and returns in 3 hours paddling at the same rate. How fast does he go in still water?
**
let x=speed in still water (mph)
c=speed of current
x+c=speed downstream
x-c=speed upstream
Distance/speed=travel time
..
6/(x+c)=40 min=2/3 hr (downstream
6/(x-c)=3 hr (upstream)
..
2x+2c=18
3x-3c=6
..
6x+6c=54
6x-6c=12
12x=66
x=66/12=33/6
ans:
speed of canoeist in still water=33/6 mph